# Monge solutions and uniqueness in multi-marginal optimal transport with hierarchical jumps

### v1

@@ 4 HJMOT on Smooth Manifolds
Here $h(\widehat{x}_{0}):=\inf\{c(\omega)\mid\omega\in\Omega,\ T_{0}(\omega)=\widehat{x}_{0}\}$ and $S(\widehat{x}_{0}):=\{\omega\in\Omega\mid T_{0}(\omega)=\widehat{x}_{0},\ c(\omega)=h(\widehat{x}_{0})\}$
@@
$D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}$ $=\lim_{n\to\infty}\frac{c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})-c(\widehat{x}_{0},\widehat{x}_{1},\cdots,\widehat{x}_{K})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}$
@@
Assume the cost function $c$ is strongly coercive and locally Lipschitz with respect to the first variable. If for $\mu_{0}$-almost every $\widehat{x}_{0}\in\mathcal{M}_{0}$, one of the following conditions holds:
  • $(1)$
    $S(\widehat{x}_{0})$ is a countable set.
  • $(2)$
    The mapping $\omega\mapsto D_{v,\,\widehat{x}_{0}}c(\omega)$ is continuous on $S(\widehat{x}_{0})$ for every $v\in T_{\widehat{x}_{0}}\mathcal{M}_{0}$.
Then the locally control condition holds.
@@
Since $c$ is locally Lipschitz continuous with respect to the initial variable $\widehat{x}_{0}$, for any $\widehat{x}_{0}\in\mathcal{M}_{0}$, there exists a Lipschitz constant $L>0$ and a neighborhood $U(\widehat{x}_{0})$ of $\widehat{x}_{0}$ such that for all $\widehat{y}_{0}\in U(\widehat{x}_{0})$ and all $(\widehat{x}_{1},\cdots,\widehat{x}_{K})\in\widehat{X}_{1}\times\cdots\times\widehat{X}_{K}$,
@@
For any $\varepsilon_{1}>0$, there exists a path $\omega^{y}=(\widehat{y}_{0},\widehat{y}_{1},\cdots,\widehat{y}_{K})\in\Omega$ such that
@@
and hence $h(\widehat{x}_{0})-h(\widehat{y}_{0})\leq\varepsilon_{1}+L\cdot{\sf d}_{0}(\widehat{x}_{0},\widehat{y}_{0})$. For any $\varepsilon_{2}>0$, there exists a path $\omega^{x}=(\widehat{x}_{0},\widehat{x}^{\prime}_{1},\cdots,\widehat{x}^{\prime}_{K})\in\Omega$ such that
@@
This shows that $h$ is Lipschitz continuous with respect $\widehat{x}_{0}\in\mathcal{M}_{0}$. Since $S(\widehat{x}_{0})$ is compact set and separable, there exists a countable dense subset $D=\{\omega^{1},\omega^{2},\cdots\}\subset S(\widehat{x}_{0})$. For each fixed $\omega^{i}=(\widehat{x}_{0},\widehat{x}_{1}^{i},\cdots,\widehat{x}_{K}^{i})\in D$, we define the function $f^{\omega^{i}}(y_{0}):=c(y_{0},\widehat{x}_{1}^{i},\cdots,\widehat{x}_{K}^{i})-h(y_{0})$ for $y_{0}\in\mathcal{M}_{0}$. For each fixed $\omega^{i}=(\widehat{x}_{0},\widehat{x}_{1}^{i},\cdots,\widehat{x}_{K}^{i})\in D$, by Rademacher’s theorem ([MR2177410, Theorem 3.1]), there exists a set $N_{\omega^{i}}\subset\mathcal{M}_{0}$ with $\mu_{0}(N_{\omega^{i}})=0$ such that for any $\widehat{x}_{0}\in\mathcal{M}_{0}\setminus N_{\omega^{i}}$, $f^{\omega^{i}}$ is differentiable at $\widehat{x}_{0}$, i.e. for any $v\in T_{\widehat{x}_{0}}\mathcal{M}_{0}$, $f^{\omega^{i}}({\rm exp}_{\widehat{x}_{0}}(tv))=f^{\omega^{i}}(\widehat{x}_{0})+D_{v,\,\widehat{x}_{0}}f^{\omega^{i}}+o(t)$. Since $f^{\omega^{i}}(y_{0})\geq 0$ for any $y_{0}\in\mathcal{M}_{0}$ and $f^{\omega^{i}}(\widehat{x}_{0})=0$, then $D_{v,\,\widehat{x}_{0}}f^{\omega^{i}}=0$. Therefore, for any $v\in T_{\widehat{x}_{0}}\mathcal{M}_{0}$,
@@
Since $D$ is dense in $S(\widehat{x}_{0})$, for any $\omega\in S(\widehat{x}_{0})$, there exists $\omega^{n}\in D$ such that $\omega^{n}\to\omega$ as $n\to\infty$. Therefore,
@@
Chapter 4 represents a concrete realization of the abstract framework from Chapter 3 within the context of smooth Riemannian manifolds. In fact, this framework admits natural extensions to more general geometric settings, including Alexandrov spaces ([MR4734965] [MR2442054]), RCD spaces ([MR3205729] [MR2984123] [MR3381131]), and sub-Riemannian manifolds ([MR2647137] [MR1867362]). Notably, the verification of the local control condition in Theorem 4.3 relies on suitable generalizations of Rademacher’s theorem, which are available in these broader contexts, thereby enabling a seamless extension of the HJMOT framework to these more general geometric spaces.

### v2

@@ 4 HJMOT on Smooth Manifolds
Here $h(\widehat{x}_{0}):=\inf\{c(\omega)\mid\omega\in{\rm supp}(\pi^{*}),\ T_{0}(\omega)=\widehat{x}_{0}\}$ and $S(\widehat{x}_{0}):=\{\omega\in{\rm supp}(\pi^{*})\mid T_{0}(\omega)=\widehat{x}_{0},\ c(\omega)=h(\widehat{x}_{0})\}$
@@
$D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}c(\omega)$ $=\lim_{n\to\infty}\frac{c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})-c(\widehat{x}_{0},\widehat{x}_{1},\cdots,\widehat{x}_{K})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}$
@@
Assume the cost function $c$ is strongly coercive and locally Lipschitz with respect to the first variable and $\mu_{0}\ll\mathcal{L}$. If for $\mu_{0}$-almost every $\widehat{x}_{0}\in\mathcal{M}_{0}$, the mapping $\omega\mapsto D_{v,\,\widehat{x}_{0}}c(\omega)$ is continuous on $S(\widehat{x}_{0})$ for every $v\in T_{\widehat{x}_{0}}\mathcal{M}_{0}$. Then the locally control condition holds.
@@
Since $c$ is locally Lipschitz continuous with respect to the initial variable $\widehat{x}_{0}$, for any $\widehat{x}_{0}\in\Omega$, there exists a Lipschitz constant $L>0$ and a neighborhood $U(\widehat{x}_{0})$ of $\widehat{x}_{0}$ such that for all $\widehat{y}_{0}\in U(\widehat{x}_{0})$ and all $(\widehat{x}_{1},\cdots,\widehat{x}_{K})\in\widehat{X}_{1}\times\cdots\times\widehat{X}_{K}$,
@@
For any $\varepsilon_{1}>0$, there exists a path $\omega^{y}=(\widehat{y}_{0},\widehat{y}_{1},\cdots,\widehat{y}_{K})\in{\rm supp}(\pi^{*})$ such that
@@
and hence $h(\widehat{x}_{0})-h(\widehat{y}_{0})\leq\varepsilon_{1}+L\cdot{\sf d}_{0}(\widehat{x}_{0},\widehat{y}_{0})$. For any $\varepsilon_{2}>0$, there exists a path $\omega^{x}=(\widehat{x}_{0},\widehat{x}^{\prime}_{1},\cdots,\widehat{x}^{\prime}_{K})\in{\rm supp}(\pi^{*})$ such that
@@
This shows that $h$ is Lipschitz continuous with respect $\widehat{x}_{0}\in\mathcal{M}_{0}$. Since ${\rm supp}(\pi^{*})$ is closed set and separable, there exists a countable dense subset $D=\{\omega^{1},\omega^{2},\cdots\}\subset{\rm supp}(\pi^{*})$. For each fixed $\omega^{i}=(\widehat{x}_{0},\widehat{x}_{1}^{i},\cdots,\widehat{x}_{K}^{i})\in D$, we define the function $f^{\omega^{i}}(y_{0}):=c(y_{0},\widehat{x}_{1}^{i},\cdots,\widehat{x}_{K}^{i})-h(y_{0})$ for $y_{0}\in\mathcal{M}_{0}$. For each fixed $\omega^{i}=(\widehat{x}_{0},\widehat{x}_{1}^{i},\cdots,\widehat{x}_{K}^{i})\in D$, by Rademacher’s theorem ([8, Theorem 3.1]), there exists a set $N_{\omega^{i}}\subset\mathcal{M}_{0}$ with $\mu_{0}(N_{\omega^{i}})=0$ such that for any $y_{0}\in\mathcal{M}_{0}\setminus N_{\omega^{i}}$, $f^{\omega^{i}}$ is differentiable at $y_{0}$, i.e. for any $v\in T_{\widehat{x}_{0}}\mathcal{M}_{0}$, $f^{\omega^{i}}({\rm exp}_{\widehat{x}_{0}}(tv))=f^{\omega^{i}}(\widehat{x}_{0})+D_{v,\,\widehat{x}_{0}}f^{\omega^{i}}+o(t)$. If $\widehat{x}_{0}\in\mathcal{M}_{0}\setminus N_{\omega^{i}}$, since $f^{\omega^{i}}(y_{0})\geq 0$ for any $y_{0}\in\mathcal{M}_{0}$ and $f^{\omega^{i}}(\widehat{x}_{0})=0$, then $D_{v,\,\widehat{x}_{0}}f^{\omega^{i}}=0$. Therefore, for any $v\in T_{\widehat{x}_{0}}\mathcal{M}_{0}$,
@@
Since $D$ is dense in ${\rm supp}(\pi^{*})$, for any $\omega\in S(\widehat{x}_{0})$, there exists $\omega^{n}\in D$ such that $\omega^{n}\to\omega$ as $n\to\infty$. Therefore,
@@
We note that, although in this chapter we have assumed all spaces to be smooth Riemannian manifolds for convenience, the proofs of the existence and uniqueness of Monge solutions actually only require the initial space $\mathcal{M}_{0}$ to be a complete smooth Riemannian manifold. The other spaces $\mathcal{M}_{1},\ldots,\mathcal{M}_{K}$ can be taken to be complete separable metric spaces (Polish spaces), as the key differential structures (differentiability along geodesics) and the twist condition are only needed for $\mathcal{M}_{0}$
@@
Chapter 4 represents a concrete realization of the abstract framework from Chapter 3 within the context of smooth Riemannian manifolds. In fact, this framework admits natural extensions to more general geometric settings, including Alexandrov spaces ([1] [3]), RCD spaces ([2] [6] [7]), and sub-Riemannian manifolds ([5] [14]). Notably, the verification of the local control condition in Theorem 4.3 relies on suitable generalizations of Rademacher’s theorem, which are available in these broader contexts, thereby enabling a seamless extension of the HJMOT framework to these more general geometric spaces.
5 The HJMOT Framework on General Non-Orientable Manifolds
Let $(\mathcal{M},g)$ be a compact, connected, smooth $d$-dimensional complete Riemannian manifold that is non-orientable. The intrinsic distance between any two points $x,y\in\mathcal{M}$ is defined as
${\sf d}_{\mathcal{M}}(x,y):=\inf\{L(\gamma)\mid\gamma\in C^{1}([0,1],\mathcal{M}),\gamma(0)=x,\gamma(1)=y\},$
where $L(\gamma)=\int_{0}^{1}\sqrt{g_{\gamma(t)}(\dot{\gamma}(t),\dot{\gamma}(t))}{\rm d}t$ is the length of curve $\gamma$.
Definition 5.1 (Double cover).
Let $\mathcal{M}$ be a connected non-orientable manifold. Its double cover consists of an orientable manifold $\widetilde{\mathcal{M}}$ and a smooth surjection $\psi:\widetilde{\mathcal{M}}\to\mathcal{M}$ satisfying:
  • $(1)$
    $\psi$ is a local diffeomorphism.
  • $(2)$
    For any $x\in\mathcal{M}$, $\psi^{-1}(x)$ contains exactly two points.
Every compact connected non-orientable manifold admits a compact connected orientable double cover (see [12]).
Definition 5.2 (Pullback Metric).
Let $(\widetilde{\mathcal{M}},\psi)$ be the double cover of $\mathcal{M}$ with Riemannian metric $g$. The pullback metric on $\widetilde{\mathcal{M}}$ is defined
$\tilde{g}:=\psi^{*}g,$
where $\psi^{*}$ denotes the pullback operation, i.e. for any point $\tilde{p}\in\widetilde{\mathcal{M}}$ and tangent vectors $u,v\in T_{\tilde{p}}\widetilde{\mathcal{M}}$, the pullback metric is given by
$\tilde{g}_{\tilde{p}}(u,v)=(\psi^{*}g)_{\tilde{p}}(u,v):=g_{\psi(\tilde{p})}({\rm d}\psi_{\tilde{p}}(u),{\rm d}\psi_{\tilde{p}}(v)).$
Here, ${\rm d}\psi_{\tilde{p}}:T_{\tilde{p}}\widetilde{\mathcal{M}}\to T_{\psi(\tilde{p})}\mathcal{M}$ is the differential of $\psi$ at $\tilde{p}$.
The pullback metric $\tilde{g}$ induces a natural Riemannian structure on $\widetilde{\mathcal{M}}$ that makes $\psi$ a local isometry. This means that in sufficiently small neighborhoods, $\psi$ preserves lengths, angles, and distances. For $\tilde{x},\tilde{y}\in\widetilde{\mathcal{M}}$, the distance on double cover is defined as
${\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x},\tilde{y}):=\inf\left.\left\{\int_{0}^{1}\sqrt{\tilde{g}_{\widetilde{\gamma}(t)}(\dot{\widetilde{\gamma}}(t),\dot{\widetilde{\gamma}}(t))}\ {\rm d}t\ \right|\ \widetilde{\gamma}\in C^{1}([0,1],\widetilde{\mathcal{M}}),\widetilde{\gamma}(0)=\tilde{x},\widetilde{\gamma}(1)=\tilde{y}\right\}$.
Lemma 5.1 (Local isometry).
The covering map $\psi:\widetilde{\mathcal{M}}\to\mathcal{M}$ is a local isometry with respect to the pullback metric $\tilde{g}=\psi^{*}g$. That is, for any $\tilde{x}\in\widetilde{\mathcal{M}}$, there exists a neighborhood $\widetilde{U}\subset\widetilde{\mathcal{M}}$ such that for all $\tilde{y},\tilde{z}\in\widetilde{U}$,
$).{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y},\tilde{z})={\sf d}_{\mathcal{M}}(\psi(\tilde{y}),\psi(\tilde{z}$)).
Proof.
For any $\tilde{x}\in\widetilde{\mathcal{M}}$, there exists a neighborhood $\widehat{U}\subset\widetilde{\mathcal{M}}$ of $\tilde{x}$ such that $\psi$ is a local diffeomorphism on $\widehat{U}$. Since $\widetilde{\mathcal{M}}$ is a Riemannian manifold, there exists a convex set $\widetilde{U}\subset\widehat{U}$. For any $\tilde{y},\tilde{z}\in\widetilde{\mathcal{M}}$,
$L_{\widetilde{\mathcal{M}}}(\widetilde{\gamma})=\int_{0}^{1}\sqrt{\tilde{g}_{\widetilde{\gamma}(t)}(\dot{\widetilde{\gamma}}(t),\dot{\widetilde{\gamma}}(t))}\ {\rm d}t=\int_{0}^{1}\sqrt{g_{\gamma(t)}({\rm d}(\psi\circ\dot{\widetilde{\gamma}})(t),\psi\circ\dot{\widetilde{\gamma}})(t))}\ {\rm d}t=L_{\mathcal{M}}(\gamma).$
[…]
$\{L_{\widetilde{\mathcal{M}}}(\widetilde{\gamma})\mid\widetilde{\gamma}\in C^{1}([0,1],\widetilde{U}),\widetilde{\gamma}(0)=\tilde{y},\widetilde{\gamma}(1)=\tilde{z}\}$
[…]
Since $\widetilde{U}$ is a convex set, the geodesic connecting $\tilde{y}$ and $\tilde{z}$ lies entirely within $\widetilde{U}$. Hence, we obtain $){\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y},\tilde{z})={\sf d}_{\mathcal{M}}(\psi(\tilde{y}),\psi(\tilde{z}$)), thus completing the proof. ∎
Lemma 5.2 (Global distance equivalence).
Let $\mathcal{M}$ be a compact connected complete non-orientable manifold with double cover $(\widetilde{\mathcal{M}},\psi)$. For any $x,y\in\mathcal{M}$, let $\psi^{-1}(x)={\tilde{x}_{1},\tilde{x}_{2}}$ and $\psi^{-1}(y)={\tilde{y}_{1},\tilde{y}_{2}}$. Then
${\sf d}_{\mathcal{M}}(x,y)=\min\{{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{i},\tilde{y}_{i})\mid i=1,2\}.$
Proof.
For any $i\in{1,2}$, let $\widetilde{\gamma}$ be a geodesic in $\widetilde{\mathcal{M}}$ connecting $\tilde{x}_{i}$ and $\tilde{y}_{j}$. Then $\gamma=\psi\circ\tilde{\gamma}$ is a curve in $\mathcal{M}$ connecting $x$ and $y$. Then,
$L_{\mathcal{M}}(\gamma)=L_{\widetilde{\mathcal{M}}}(\widetilde{\gamma})={\sf d}_{\widetilde{\mathcal{M}}}(x_{i},y_{j}).$
Hence ${\sf d}_{\mathcal{M}}(x,y)\leq{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{i},\tilde{y}_{i})$ for each $i$, then
${\sf d}_{\mathcal{M}}(x,y)\leq\min\{{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{i},\tilde{y}_{i})\mid i=1,2\}.$
Now we prove the reverse inequality. Let $\gamma$ be a geodesic in $\mathcal{M}$ connecting $x$ and $y$. Since $\psi$ is a covering map, there exists a lift $\tilde{\gamma}$ of $\gamma$ in $\widetilde{\mathcal{M}}$ such that $\psi\circ\tilde{\gamma}=\gamma$ and $\tilde{\gamma}$ connects $\tilde{x}_{i}$ to $\tilde{y}_{i}$ for $i=1,2$. Then,
${\sf d}_{\mathcal{M}}(x,y)=L_{\mathcal{M}}(\gamma)=L_{\widetilde{\mathcal{M}}}(\widetilde{\gamma})\geq{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{i},\tilde{y}_{j})\geq\min\{{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{i},\tilde{y}_{j})\mid i,j=1,2\}.$
Combining both inequalities gives the desired equality. ∎
Let $(\widetilde{\mathcal{M}},\psi)$ be the double cover of $\mathcal{M}$. Define the extended intermediate space as $\widehat{\mathcal{M}}$$\widetilde{\mathcal{M}}\cup\{\partial\}$ where $\partial$ is topological isolated point. Define the path space as $\Omega:=\mathcal{M}\times\widehat{\mathcal{M}}\times\mathcal{M}$. For a path $\omega=(x_{0},x_{1},x_{2})\in\Omega$, define the cost function as
$c(\omega):=\left\{\begin{aligned} &\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}+\inf_{\tilde{x}_{2}\in\psi^{-1}(x_{2})}{\sf d}_{\widetilde{\mathcal{M}}}(x_{1},\tilde{x}_{2})^{2},&\quad&\text{if}\quad&x_{1}\in\widetilde{\mathcal{M}},\\ &{\sf d}_{\mathcal{M}}(x_{0},x_{2})^{2},&&\text{if}&x_{1}=\partial.\end{aligned}\right$.
Theorem 5.1.
For fix $\mu_{0}\in\mathcal{P}(\mathcal{M})$ and $\mu_{2}\in\mathcal{P}(\mathcal{M})$. We seek to solve the following optimization problem
$M:=\inf_{\mu_{1}\in\mathcal{P}(\widehat{\mathcal{M}}),\,\pi\in\Pi(\mu_{0},\mu_{1},\mu_{2})}\int_{\widetilde{\Omega}}c(\omega)\ \pi({\rm d}\omega).$ (5.1)
where $\widetilde{\Omega}:=\Omega\setminus\{\omega\in\Omega\mid\psi(x_{1})=x_{0}\text{ or }\psi(x_{1})=x_{2}\}$. If $M<\infty$ and $(x_{1},x_{2})\mapsto D_{v,\,x_{0}}c(x_{0},x_{1},x_{2})$ is injective on $\widetilde{\Omega}$ for every $v\in T_{x_{0}}\mathcal{M}$ and $\mu_{0}$-a.e. $x_{0}$, then there exists a unique Borel map $T^{*}:\mathcal{M}\to\widetilde{\Omega}$ such that
  • $\rm(1)$}
    The coupling $\pi^{*}={T^{*}}_{\#}\mu_{0}$ is the unique solution of (5.1).
  • $\rm(2)$}
    For each $x_{0}\in\mathcal{M}$, $T^{*}$ decomposes as $T^{*}(x_{0})=(x_{0},T_{1}^{*}(x_{0}),T_{2}^{*}(x_{0}))$.
Proof.
We first note that the optimization problem (5.1) can be restricted without loss of generality to the subset
$\widetilde{\Omega}=\Omega\setminus\{\omega\in\Omega\mid\psi(x_{1})=x_{0}\text{ or }\psi(x_{1})=x_{2}\}.$
Indeed, for any path $\omega=(x_{0},x_{1},x_{2})\in\Omega$ with $\psi(x_{1})=x_{0}$ or $\psi(x_{1})=x_{2}$, consider the alternative path $\omega^{\prime}=(x_{0},\partial,x_{2})$. From the definition of the cost and Lemma 5.2, one verifies that $c(\omega^{\prime})\leq c(\omega)$. Consequently, for any coupling $\pi\in\Pi(\mu_{0},\mu_{1},\mu_{2})$, the mass supported on $\Omega\setminus\widetilde{\Omega}$ can be transferred to paths of the form $\omega^{\prime}$ without increasing the total cost and while preserving all marginals. Hence, there exists an optimal solution $\pi^{*}$ of (5.1) satisfying $\pi^{*}(\widetilde{\Omega})=1$. Therefore, we may and shall restrict the problem to the path space $\widetilde{\Omega}$ without affecting the optimal value or the existence of a minimizer.
Next, we verify the hypotheses of Theorem 4.2 on $\Omega$. Since $\widetilde{\Omega}$ is a subset of $\Omega$, these conditions naturally hold on $\widetilde{\Omega}$ as well. Since $\mathcal{M}$ is compact, it is clear that $c$ is strongly coercive. When $x_{1}=\partial$, since $\sf d$$_{\mathcal{M}}$ is a Riemannian distance function, it is evident that $c$ is locally Lipschitz with respect to the first variable. If $x_{1}\in\widetilde{\mathcal{M}}$, for $x_{0}\in\mathcal{M}$, there exists a neighborhood $U\subset\mathcal{M}$ of $x_{0}$ such that for any $y_{0}\in U$, there exists $\tilde{x}_{0}^{1},\tilde{x}_{0}^{2}\in\psi^{-1}(x_{0})$ and $\tilde{y}_{0}^{1},\tilde{y}_{0}^{2}\in\psi^{-1}(y_{0})$, we have
${\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},\tilde{y}_{0}^{i})={\sf d}_{\mathcal{M}}(x_{0},y_{0})\quad\text{for}\quad i=1,2.$
By the Lipschitz continuity of the squared distance function on compact manifolds, there exists $L>0$ such that for $i=1,2$,
$\left|{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},x_{1})^{2}-{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y}_{0}^{i},x_{1})^{2}\right|\leq L\cdot{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},\tilde{y}_{0}^{i})=L\cdot{\sf d}_{\mathcal{M}}(x_{0},y_{0}).$
Set $k=\arg\min_{i=1,2}{{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},x_{1})}$, then
$\min_{i=1,2}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y}_{0}^{i},x_{1})^{2}\leq{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y}_{0}^{k},x_{1})^{2}\leq{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{k},x_{i})^{2}+L\cdot{\sf d}_{\mathcal{M}}(x_{0},y_{0})=\min_{i=1,2}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},x_{1})^{2}+L\cdot{\sf d}_{\mathcal{M}}(x_{0},y_{0}).$
Similarly, set $m=\arg\min_{i=1,2}{{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y}_{0}^{i},x_{1})}$, we have
$\min_{i=1,2}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},x_{1})^{2}\leq{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{m},x_{1})^{2}\leq{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y}_{0}^{m},x_{i})^{2}+L\cdot{\sf d}_{\mathcal{M}}(x_{0},y_{0})=\min_{i=1,2}{{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y}_{0}^{i},x_{1})}^{2}+L\cdot{\sf d}_{\mathcal{M}}(x_{0},y_{0}).$
Combining the two inequalities above, we obtain for $x_{0}\in\mathcal{M}$ , there exists a neighborhood $U$ of $x_{0}$ such that
$\left|\min_{i=1,2}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},x_{1})^{2}-\min_{i=1,2}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{y}_{0}^{i},x_{1})^{2}\right|\leq L\cdot{\sf d}_{\mathcal{M}}(x_{0},y_{0}).$
Thus the cost function $c$ is locally Lipschitz with respect to the first variable. When $x_{1}=\partial$, since $\partial$ is an isolated point, for any $(x_{1}^{n},x_{2}^{n})$ converging to $(x_{1},x_{2})$, when $n$ is sufficiently large, $x_{1}^{n}=\partial$. Hence $c(x_{0},x_{1}^{n},x_{2}^{n})={\sf d}_{\mathcal{M}}(x_{0},x_{2}^{n})^{2}$. Since the Riemannian metric is smooth function on a smooth manifold, $(x_{1},x_{2})\mapsto D_{v,\,x_{0}}c(x_{0},x_{1},x_{2})$ is continuous. When $x_{1}\in\widetilde{\mathcal{M}}$, set $f_{i}(x_{0},x_{1}):={\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0}^{i},x_{1})^{2}$ for $i=1,2$. If $f_{1}\neq f_{2}$, since $f_{1}$ and $f_{2}$ is continuous on $\mathcal{M}\times\widetilde{\mathcal{M}}$, then there exists a neighborhood $U\subset\mathcal{M}\times\widetilde{\mathcal{M}}$ of $(x_{0},x_{1})$ such that the order relation between $f_{1}$ and $f_{2}$ remains unchanged within $U$. Therefore, for any $(x_{0},x_{1})\in U$, if $f_{1}<f_{2}$, then $\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}=f_{1}(x_{0},x_{1})$. If $f_{1}>f_{2}$, then $\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}=f_{2}(x_{0},x_{1})$. Since $f_{1}$ and $f_{2}$ are smooth at $\mu_{0}$-a.e. $x_{0}$, $\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}$ is also smooth in $U$, thus $D_{v,\,x_{0}}\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}$ is continuous. If $f_{1}(x_{0},x_{1})=f_{2}(x_{0},x_{1})$, we have
$\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}=f_{1}(x_{0},x_{1})=f_{2}(x_{0},x_{1}).$
For $\mu_{0}$-a.e. $x_{0}$, we have $D_{v,\,x_{0}}f_{1}=D_{v,\,x_{0}}f_{2}$ for any $v\in T_{x_{0}}\mathcal{M}$. Therefore $D_{v,\,x_{0}}\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}$ $D_{v,\,x_{0}}f_{1}(x_{0},x_{1})=D_{v,\,x_{0}}f_{2}(x_{0},x_{1})$. Then $D_{v,\,x_{0}}\inf_{\tilde{x}_{0}\in\psi^{-1}(x_{0})}{\sf d}_{\widetilde{\mathcal{M}}}(\tilde{x}_{0},x_{1})^{2}$ is continuous in $(x_{0},x_{1})$ for $\mu_{0}$-a.e. $x_{0}$. Finally, we have $\omega\mapsto D_{v,\,x_{0}}c(\omega)$ is continuous on $\widetilde{\Omega}$ for every $v\in T_{x_{0}}\mathcal{M}$. According to Theorem 4.3, we know that $c$ satisfies the locally control condition. Finally, according to Theorem 4.2, we complete the proof. ∎
If for $\mu_{0}$-a.e. $x_{0}\in\mathcal{M}$, the optimal intermediate point satisfies $T_{1}^{*}(x_{0})=\partial$, then the HJMOT problem reduces to the classical two-marginal optimal transport between $\mu_{0}$ and $\mu_{2}$ with cost function $c(x_{0},x_{2})={\sf d}_{\mathcal{M}}(x_{0},x_{2})^{2}$. In this case, the map $T_{2}^{*}:\mathcal{M}\to\mathcal{M}$ becomes the unique Monge solution of the classical OT problem.
The HJMOT framework demonstrates a powerful strategy for extending optimal transport theory to challenging geometric settings. By introducing well-behaved intermediate spaces (in this case, the orientable double cover $\widetilde{\mathcal{M}}$) and allowing mass to jump selectively through these spaces, we effectively restrict the computational complexity to these ”nice” spaces. For non-orientable manifolds $\mathcal{M}$, the problematic geometric features are circumvented by transporting through the orientable double cover $\widetilde{\mathcal{M}}$ when beneficial, while still permitting direct transport when the intermediate space offers no advantage. This approach enables us to leverage the rich theory of optimal transport on smooth orientable manifolds to solve problems on more challenging spaces, thus significantly extending the applicability of Monge solutions beyond classical settings.